# Protein Conformation Generation via Force-Guided $\mathrm{SE}(3)$ Diffusion Models

### v1

@@ C.4 Details on evaluation metrics
where $KL(\cdot\|\cdot)$ is the Kullback-Leibler divergence. Similar to (Lu et al., 2024), we project coordinates to three types of feature spaces to compute conformation distribution: 1) Pairwise Distance (PwD) between $\alpha$-carbon of residues excluding adjacent residues within an offset of 3; 2) Radius of gyrations (Rg) that are distances of $\alpha$-carbon atoms to the center-of-mass of the conformation; 3) Time-lagged Independent Components (TIC) from TICA analysis on the reference MD trajectories (see Appendix C.2), only the first two components are used for JS calculation. Feature dimensions are discretized with 50 bins according to the values of the reference MD samples and a pseudo count of 1e-6 is added. We compute the distance over each dimension and report the mean values as the final JS distance.
@@ D.1 Ablation studies on force and energy guidance
Figure S2: Energy (left) and diversity (RMSF, right) of sampled conformation for WW doamin with various levels of energy guidance ($\eta$) and sequence conditions ($\gamma$).
@@
D.2 Sample distribution over TIC components for all 12 fast-folding proteins
Here we include sample distributions over the first two dimension of TIC components for all 12 fast-folding proteins.
Figure S3: Sample distribution over the first two TIC components for all 12 fast-folding proteins. Counters indicate the full reference MD sample density at log-scale.
D.3 Additional results for BPTI
We list the best RMSD scores to each 5 cluster centers in Table S3. ConfDiff-Force and EigenFold can most accurately predict the structures of metastable states. Especially, ConfDiff-Force out performs all other models in predicting the remote cluster 3.
Table S3: Detailed best RMSD to each cluster center. Unit: $\mathrm{\SIUnitSymbolAngstrom}$
@@ D.3 Additional results for BPTI
Figure S4: Model generated BPTI conformations projected to the first two TIC. Contour lines represent the ground truth probability density at the log scale.
D.4 ConfDiff-Base results on Apo-holo
Apo-holo is a dataset consists of 90 proteins with known conformation changes upon ligand-binding and each has two conformations (apo, the unbound state; holo, the bound state) determined by experiments. This dataset introduces a challenging task to predicting distinct conformations with various protein sizes and structural changes (Jing et al., 2023; Saldaño et al., 2022). Due to the high computation cost of evaluating the energy and force for mid-to-large size structure using OpenMM, we only compare the performance of our classifier-free model with baselines.
@@ D.4 ConfDiff-Base results on Apo-holo
[…]
@@
Table S4: Apo-holo results (90 casee). TM-Div is the mean pairwise TMscore between samples.

### v2

@@ C.4 Details on Evaluation Metrics
where $\mathrm{KL}(\cdot\|\cdot)$ is the Kullback-Leibler divergence. Similar to (Lu et al., 2024), we project coordinates to three types of feature spaces to compute conformation distribution: 1) Pairwise Distance (PwD) between $\alpha$-carbon of residues excluding adjacent residues within an offset of 3; 2) Radius of gyrations (Rg) that are distances of $\alpha$-carbon atoms to the center-of-mass of the conformation; 3) Time-lagged Independent Components (TIC) from TICA analysis on the reference MD trajectories (see Appendix C.2), only the first two components are used for JS calculation. Feature dimensions are discretized with 50 bins according to the values of the reference MD samples and a pseudo count of $10^{-6}$ is added. We compute the distance over each dimension and report the mean values as the final JS distance.
@@ D.1 Ablation Studies on Force and Energy Guidance
Figure S2: Energy (left) and diversity (RMSF, right) of sampled conformation for WW domain with various levels of energy guidance ($\eta$) and sequence condition ($\gamma$).
@@
D.2 Additional Discussion on Fast-folding Results
Figure S3: Five worst structures with highest residue clash rate. Both the cartoon representation and backbone bonds are shown to better depict the clashes.
We find that Jensen-Shannon (JS) metrics alone are insufficient to reflect the quality and diversity of sampled conformations. Note that the JS distance here is calculated and averaged over each dimension (e.g., $\alpha$-carbons), which is not able to capture the bond breaking and steric clash. With the introduction of energy/force guidance, irrational conformations are rectified and structures tend to contract towards the local minimum. This is the main factor leading to a performance drop in distributional metrics.
Empirically, ConfDiff-base sometimes samples poor structures with steric clashes, while energy/force guidance strategies enable sampling high-fidelity conformations (i.e., high VAL-CA scores) with slightly worse JS metrics. As shown in Figure 2 in our paper, RMSF (reflecting structural diversity) values fluctuate with various force guidance strengths, while the potential energy keeps improving. This also explains the fluctuation of JS metrics with improved VAL-CA scores after energy/force guidance.
Here we show some examples to compare sampled structures from ConfDiff-base and ConfDiff-force. Among 1000 samples from ConfDiff-base ($\gamma=0.5$) and ConfDiff-force ($\gamma=0.5$, $\eta=1.3$), we select five worst conformations (with highest clash rate) shown in the Figure S3. ConfDiff-force samples exhibit fewer steric clashes, demonstrating more realistic protein structures, despite slightly worse JS distance. Therefore, it is more appropriate to take into account all metrics for a comprehensive assessment of the generated conformation ensemble.
D.3 Sample Distribution over TIC Components for All Fast-Folding Proteins
Here we include sample distributions over the first two TIC components for all twelve fast-folding proteins.
Figure S4: Sample distribution over the first two TIC components for twelve fast-folding proteins. Contours indicate the full reference MD sample density at log-scale.
D.4 Additional Results for BPTI
We list the best RMSD scores to each five cluster centers in Table S5. ConfDiff-Force and EigenFold can most accurately predict the structures of metastable states. Especially, ConfDiff-Force out-performs all other models in predicting the remote cluster 3.
Table S5: Detailed best RMSD to each cluster center. Unit: $\mathrm{\SIUnitSymbolAngstrom}$.
@@ D.4 Additional Results for BPTI
Figure S5: Model generated BPTI conformations projected to the first two TIC components. Contour lines represent the ground truth probability density at the log scale.
D.5 ConfDiff-Base Results on Apo-holo Dataset
Apo-holo is a dataset consists of 90 proteins with known conformation changes upon ligand-binding and each has two conformations (apo, the unbound state; holo, the bound state) determined by experiments. This dataset introduces a challenging task of predicting distinct conformations of proteins with various sizes and structural changes (Jing et al., 2023; Saldaño et al., 2022). Due to the high computation cost of evaluating the energy and force for mid-to-large size structure using OpenMM, we only compare the performance of our classifier-free model with baselines.
@@ D.5 ConfDiff-Base Results on Apo-holo Dataset
$\textsc{TMens}=1/2\times\Big{(}\max_{i}\big{\{}\text{TMscore}(\text{sample }i,\textit{apo})\big{\}}+\max_{i}\big{\{}\text{TMscore}(\text{sample }i,\textit{holo})\big{\}}\Big{)}$
@@
Table S6: Apo-holo results (90 cases). TM-Div is the mean pairwise TMscore between samples.
@@
D.6 ConfDiff Results on Sars-CoV-2
In order to fully understand the method’s generalization capabilities, we did some proof-of-concept experiments to extend proposed method to challenging cases of large proteins - the SARS-CoV-2 case studies introduced in Distributional Graphormers (DiG) (Zheng et al., 2023). This dataset contains two important proteins in drug discovery: the spike receptor-binding proteins (RBD, 194 amino acids) and the main protease (Mpro, 306 amino acids). MD simulations for both proteins are publicly available and are used as reference during evaluation.
Preliminary results showed that protein sequence information using representations from ESMFold (Lin et al., 2022) is not able to predict reasonable structures of RBD and Mpro, while AlphaFold predictions are close to the experimental structures in Table S7.
Table S7: The RMSD and TMscore of RBD and Mpro from different sequence representations.
RMSD TMscore
ESMFold_RBD 20.55 0.18
ESMFold_Mpro 13.96 0.59
AlphaFold_RBD 2.55 0.91
AlphaFold_Mpro 3.28 0.93
For quantitative comparison, we calculated the following metrics: 1) average RMSD to experimental structures (PDB IDs are RBD: 6LU7, Mpro: 6M0J); 2) the JS-TIC2D to reference distribution; 3) performances on the binary classification of TICA states similar to DiG: the joint space is discretized into $50\times 50$ states and those with/without ground truth conformations are positive/nagtive states; 4) diversity as pairwise RMSD between samples (RMSF) We summarized the metrics below in Table S8, with some major findings:
Table S8: Results on Sars-CoV-2. The results for classifier-free, energy guidance, and force guidance are labeled as ConfDiff-Base, ConfDiff-Force, and ConfDiff-Energy. All values are shown as mean.
Binary prediction of TIC2D states
Model RMSD (Å) JS-TIC2D $\downarrow$ Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ RMSF (Å)
RBD
ConfDiff-Base 1.396 0.822 0.72 0.147 0.244 1.55
ConfDiff-Energy 1.341 0.826 0.647 0.133 0.221 1.564
ConfDiff-Force 1.385 0.825 0.765 0.154 0.257 1.606
Mpro
ConfDiff-Base 1.475 0.785 1 0.125 0.221 1.692
ConfDiff-Energy 1.448 0.794 1 0.116 0.208 1.722
ConfDiff-Force 1.559 0.768 1 0.132 0.234 1.828
  • All the models can accurately predict the structure for both RBD and Mpro with mean RMSD$<$1.6 Å. The samples are scattered near the major peaks in the reference distribution as a single cluster. For RBD, the samples spread out in the same direction as the reference distributions.
  • Adding energy or force guidance does not significantly impact the metrics. While ConfDiff-Force showed some positive signs on slightly improved metrics (e.g., JS-TIC2D for Mpro and Recall & F1 scores for both proteins).
  • The current prototype is only fine-tuned on the AF2 representation with two proteins (RBD and Mpro). Lack of training on large datasets could significantly impact model’s performance. We could train the base model on the entire PDB structures with AF2 representations in the future and we believe the model’s performance would greatly improve.
  • Nevertheless, we aim to use these experiments to show that 1) proposed methods can scale up and apply to large proteins with real-world applications; 2) even with much simplification on experiments, the results showed some promising signs on proposed results.